find four consecutive even integers whose sum is 340
Answer by feliz1965(151)

Here are your 4 consecutive even integers:
x
x + 2
x + 4
x + 6
=========================================
We now add them and equate the whole thing to 340. Then we solve for x.
By the way, you can use letter. I always use x, y or z because they are popular letters in algebra.
x + x + 2 + x + 4 + x + 6 = 340
4x + 12 = 340
4x = 340 - 12
4x = 328
x = 328/4
x = 82
=========================================
Now that we know x, we plug that value of x into the 4 consecutive numbers given above and simplify (where needed).
Here are the 4 numbers:
82, 84, 86, 88